A physics student spends part of her day walking between classes or for recreation, during which time she expends energy at an average rate of 290 W . The remainder of the day she is sitting in class, studying, or resting; during these activities, she expends energy at an average rate of 120 W .

If she expends a total of 1.2×107 J of energy in a 24-hour day, how much of the day did she spend walking?
Answer #1

Pavg = 1.2 x10^7/(24*3600s) = 138.89 W

Pavg = (290*x + 120 * (24-x) )/24

Pavg = (170x + 120*24)/24

P*24 = 170x + 120*24

170x = 24 ( 138.89 - 120)/170

x = 2.67 h

walkjin 24 - 2.67 = 21.33 hrs